OpenStudy (anonymous):

Choose a number between 55 and 101 that is a multiple of 4 , 6 , and 8. write all the numbers that be choosen

OpenStudy (zzr0ck3r):

4*6*8 = ? does this satisfy the conditions?

OpenStudy (anonymous):

is this the correct answer? because im having trouble with multiple factors problems

OpenStudy (zzr0ck3r):

what is \[4\times 6\times 8\]?

OpenStudy (anonymous):

is it 192?

OpenStudy (zzr0ck3r):

yes, does this satify the condition of being less than 101 and greater than 55?

OpenStudy (anonymous):

I dont think so? plus is it the only factor?

OpenStudy (zzr0ck3r):

well then any multiple of that wont fall within the range we need so there are no numbers that satisfy what the question is asking

OpenStudy (anonymous):

oh okay well thank you :)